# Optimizing Learning for Robust Hyperbolic Deep Learning in Computer Vision

### v2

@@ 3.2 Riemannian Optimization
Algorithm 1 Curvature Learning Aware Optimization Given parameters $\bm{\theta}=[\bm{\theta}_{euclid},\bm{\theta}_{\mathbb{L}},\bm{\theta}_{K}]$ and gradients denoted $\mathcal{G}$ procedure Optimizer Step($\bm{\theta}$)      for $\bm{p}$ in $\bm{\theta}_{\mathbb{L}}$ do          Vanilla Riemannian Optimizer Step      end for      for $\bm{p}$ in $\bm{\theta}_{euclid}\cup\bm{\theta}_{K}$ do          Vanilla Eucldiean Optimizer Step      end for      MoveParameters() end procedure procedure Move Parameters      for $\bm{p}$ in $\bm{\theta}_{\mathbb{L}}$ do          $\bm{\mathcal{G}}_{temp}=\mathrm{PT}^{K_{t-1}}_{{\bm{p}}\rightarrow{\bm{\overline{\bm{0}}}_{t-1}}}\left({\bm{\mathcal{G}}}\right)$          $\bm{z}=\log_{\bm{\overline{\bm{0}}}_{t-1}}^{K_{t-1}}(\bm{p})$          $\bm{p}=\exp_{\overline{\bm{0}}_{t}}^{K_{t}}(\bm{z})$          $\bm{\mathcal{G}}=\mathrm{PT}^{K_{t}}_{{\bm{\overline{\bm{0}}}_{t}}\rightarrow{\bm{p}}}\left({\bm{\mathcal{G}}_{temp}}\right)$      end for end procedure
Maximum Distance Rescaling
@@
To constrain hyperbolic vectors within a specified maximum distance, either a normalization function or a parameter clipping method is required. Parameter clipping can be challenging to train, as it may lead to information loss and introduce non-smooth gradients. On the other hand, common normalization functions like tanh and the sigmoid function tend to saturate quickly, limiting their effectiveness as seen in the sigmoid implementation by Chen et al. (2022). To address these issues, we introduce a modified scaling function, designed to provide finer control over both the maximum values and the slope of the curve. A visualization of this function is provided in Figure 3, and the formulation is presented below:
@@
where $y\in\mathbb{R}^{d}$, $m$ is our desired maximum value, and $s$ controls the slope of the curve. We now have a maximum distance value to adhere to and a flexible distance normalizing function. To apply this to the hyperbolic vectors, we suggest performing the scaling on the tangent plane of the origin. However, this is an expensive operation to perform often, as such we derive in appendix A.2 the equivalent factorized form for the scaling of the space values:
@@
where $D(\bm{x},\overline{\bm{0}})_{rescaled}^{k}$ is the scaled distances by plugging in $D(\bm{x},\overline{\bm{0}})$ and $D^{K}_{\overline{\bm{0}}_{max}}$ in Eq.2. The time component of the vector is then recalculated based on the norm of its spatial components. This gives us a complete scaling operation for the Lorentz space. We apply this tanh scaling when moving parameters across different manifolds. This includes transitions from the Euclidean space to the Lorentz space, as well as between Lorentz spaces of different curvatures. Additionally, we use this scaling after Lorentz Boosts and direct Lorentz concatenations (Qu & Zou, 2022). We also incorporate it following the variance-based rescaling in the batch normalization layer, since variance adjustments during this operation can push points outside the radius during the operation and often lead to invalid values.
3.3 Towards Efficient Architectural Components
Lorentz Convolutional Layer
In their work, Bdeir et al. (2024) proposed a fully hyperbolic 2D convolutional layer by breaking down the convolution operation into a window-unfolding step followed by a modified version of the Lorentz Linear Layer from Chen et al. (2022). This approach ensured that the convolution outputs remained on the hyperboloid. However, the manual patch creation combined with matrix multiplication made the computation extremely expensive, as it prevented the use of highly optimized CUDA implementations for convolutions.
To address this issue, we adopt an alternative definition of the Lorentz Linear layer from Dai et al. (2021), which decomposes the transformation into a Lorentz boost and a Lorentz rotation. Using this definition, we replace the matrix multiplication employed by Bdeir et al. (2024) for the spatial dimensions and time component projection with a learned rotation operation and a Lorentz boost. Additionally, we can achieve the rotation operation using a parameterization of the convolution weights while still relying on the CUDA convolution implementations, significantly improving computational efficiency.
A rotation is any operation that modifies the directions of transformed vectors while preserving their magnitudes. This can be achieved using a matrix $\bm{\hat{W}}\in\mathbb{S}$. Where
$\mathbb{S}(n^{\prime},n)={\bm{M}\in\mathbb{R}^{(n^{\prime}\cdot n)}:\bm{M}^{T}\bm{M}=\bm{I}}$ (4)
represents the Stiefel manifold. The condition $n\leq n^{\prime}$ ensures that the columns of the weight matrix are orthogonal, thereby preserving the norms along the rows of the input matrix, which correspond to the individual embeddings in this context.
To apply this concept to the convolution operation, the convolution weights, after unfolding, must form a rotation matrix. We define the dimensions of this matrix as $n=(channels_{in}\cdot kernel_{width}\cdot kernel_{height})$ and $n^{\prime}=channels_{out}$ respectively. However, the condition for orthogonal columns is not always satisfied as there will be instances where $n>n^{\prime}$. For these cases we use the norm-preserving transformation $\bm{z}=\bm{W^{T}}\bm{x}\cdot\frac{\lVert\bm{x}\rVert}{\lVert\bm{W^{T}}\bm{x}\rVert}$ where $\bm{W}\in\mathbb{R}^{(n^{\prime}\cdot n)}$. This formulation allows us to utilize existing efficient implementations of the convolution operation by directly parameterizing the kernel weights before passing them into the convolutional layer. Finally, we formalize the new Lorentz Convolution as:
$\bm{out}=\text{LorentzBoost(TanhScaling(RotationConvolution(}\bm{x}\textbf{)))}$)) (5)
where TanhRescaling is the operation described in Eq.3 and RotationConvolution is a normal convolution parameterized through the procedure in Algorithm 2, where Orthogonalize is a Cayley projection similar to (Helfrich et al., 2018) or the norm-preserving transformation above. We specifically use the Cayley projection because it always produces an orthonormal matrix with a positive determinant, ensuring that the rotated point remains on the upper sheet of the hyperboloid and avoids being mapped to the lower sheet.
Algorithm 2 Lorentz Convolution Parametarization
$\bm{W}\in\mathbb{R}^{C_{in},C_{out},K_{width},K_{length}}$
procedure Adaptweight(W)
     if $K_{width}\cdot K_{length}\cdot C_{in}>C_{out}$ then
         return $\bm{W}$
     else
         $\bm{W}=\bm{W}.reshape(K_{width}\cdot K_{length}\cdot C_{in},C_{out})$
         $\bm{\hat{W}}_{core}=$ Orthogonalize ($\bm{W}$)
         return $\bm{\hat{W}}.reshape(C_{in},C_{out},K_{width},K_{length})$
     end if
end procedure
$\bm{\hat{W}}_{core}$ can also be learned directly on the SPD manifold similar to Dai et al. (2021). This definition of the convolution operation allows to use the existing efficient implementations of 2D convolutions, saving both memory and runtime.
Lorentz-Core Bottleneck Block
To build on the concept of hybrid hyperbolic encoders introduced by (Bdeir et al., 2024), we developed the Lorentz Core Bottleneck blocks for hyperbolic ResNet-based models. These blocks are similar to standard Euclidean bottleneck blocks but replace the internal 3x3 convolutional layer with our efficient convolutional layer as illustrated in figure 2. This design allows us to incorporate hyperbolic structuring of the embeddings within each block while retaining the flexibility and computational efficiency of Euclidean models. We interpret this integration as a form of hyperbolic bias, enabling ResNets to leverage hyperbolic representations without requiring fully hyperbolic modeling.
Euclidean 1x1 ConvManifold Switcher + ScaleHyperbolic 3x3 ConvManifold Switcher + ScaleEuclidean 1x1 ConvBN + ReLUHyper BN + ReLU
Figure 2: Lorentz-Core Bottleneck Block
4 Experiments
@@ 4.1 Hierarchical Metric Learning Problem
We follow the experimental setup in Kim et al. (2023) and rely on four main datasets: CUB-200-2011 (CUB)(Welinder et al., 2010), Cars-196 (Cars)(Krause et al., 2013), Stanford Online Product (SOP)(Song et al., 2016), and In-shop Clothes Retrieval (InShop)(Liu et al., 2016). Performance is measured using Recall@k which is the fraction of queries with one or more relevant samples in their k-nearest neighbors. Additionally, all model backbones are pre-trained on Imagenet to ensure fair comparisons with the previous work.
@@
To adapt the HIER model to the hyperboloid we first replace the Euclidean linear layer with a Lorentzian linear layer in the model neck and implement our max distance scaling operation after. We then set the hierarchical proxies as learnable hyperbolic parameters and optimize them directly on the manifold using our Lorentzian AdamW. Finally, we change the Poincaré distance to the Lorentz distance for the LHIER loss and set the hierarchical proxies to be scaled beforehand. We continue to use FP16 precision during curvature learning to evaluate the stability of the new optimization scheme. Our experiments also include both CNN-based and transformer models to assess whether differences in architecture affect the performance of the hyperbolic heads.
@@ Results
As shown in table 1, our HIER+ model manages to improve the performance of mainly the CNN-based models with improvements ranging in recall@1 from 1-2.3%. In contrast, transformer models exhibit marginal gains in most cases and even show slight declines in other scenarios. This indicates that the proposed components do not better fully exploit the advantages of transformer architectures compared to other hyperbolic methods. The reasons may include architectural limitations or insufficient hyperparameter tuning. Notably, Kim et al. (2023) employed an extensive set of hyperparameters for model training, which we kept at default values to ensure fair comparisons and avoid costly hyperparameter searches.
Table 1: Performance of metric learning methods on the four datasets as provided by (Kim et al., 2023). All architecture backbones are pretrained and tested with the new LHIER loss. Superscripts denote their embedding dimensions and $\dagger$ indicates models using larger input images. As in (Kim et al., 2023), network architectures are abbreviated as, R–ResNet50 (He et al., 2016), B–Inception with BatchNorm (Ioffe & Szegedy, 2015), De–DeiT (Touvron et al., 2021), DN–DINO (Caron et al., 2021) and V–ViT (Dosovitskiy et al., 2021)
Methods Arch. CUB Cars SOP
R@1 R@2 R@4 R@1 R@2 R@4 R@1 R@10 R@100
Backbone architecture: CNN
NSoftmax (Zhai & Wu, 2018) R128 56.5 69.6 79.9 81.6 88.7 93.4 75.2 88.7 95.2
MIC (Roth et al., 2019) R128 66.1 76.8 85.6 82.6 89.1 93.2 77.2 89.4 94.6
XBM (Wang et al., 2020) R128 - - - - - - 80.6 91.6 96.2
XBM (Wang et al., 2020) B512 65.8 75.9 84.0 82.0 88.7 93.1 79.5 90.8 96.1
HTL (Ge et al., 2018) B512 57.1 68.8 78.7 81.4 88.0 92.7 74.8 88.3 94.8
MS (Wang et al., 2019) B512 65.7 77.0 86.3 84.1 90.4 94.0 78.2 90.5 96.0
SoftTriple (Qian et al., 2019) B512 65.4 76.4 84.5 84.5 90.7 94.5 78.6 86.6 91.8
PA (Kim et al., 2020) B512 68.4 79.2 86.8 86.1 91.7 95.0 79.1 90.8 96.2
NSoftmax (Zhai & Wu, 2018) R512 61.3 73.9 83.5 84.2 90.4 94.4 78.2 90.6 96.2
ProxyNCA++ (Teh et al., 2020) R512 69.0 79.8 87.3 86.5 92.5 95.7 80.7 92.0 96.7
Hyp (Ermolov et al., 2022) R512 65.5 76.2 84.9 81.9 88.8 93.1 79.9 91.5 96.5
HIER (Kim et al., 2023) R512 70.1 79.4 86.9 88.2 93.0 95.6 80.2 91.5 96.6
LHIER (ours) R512 72.4 81.5 88.4 89.1 93.5 96.1 81.3 92.1 96.8
Backbone architecture: ViT
[…]
Hyp (Ermolov et al., 2022) De128 74.7 84.5 90.1 82.1 89.1 93.4 83.0 93.4 97.5
HIER (Kim et al., 2023) De128 75.2 84.2 90.0 85.1 91.1 95.1 82.5 92.7 97.0
LHIER (ours) De128 75.5 84.7 90.6 85.4 91.7 95.6 82.7 93.5 97.4
Hyp (Ermolov et al., 2022) DN128 78.3 86.0 91.2 86.0 91.9 95.2 84.6 94.1 97.7
HIER (Kim et al., 2023) DN128 78.5 86.7 91.5 88.4 93.3 95.9 84.9 94.2 97.5
LHIER (ours) DN128 78.8 87.0 91.9 88.9 93.2 96.4 85.1 94.9 98.2
Hyp (Ermolov et al., 2022) V128 84.0 90.2 94.2 82.7 89.7 93.9 85.5 94.9 98.1
HIER  (Kim et al., 2023) V128 84.2 90.1 93.7 86.4 91.9 95.1 85.6 94.6 97.8
LHIER (ours) V128 84.6 90.2 93.9 86.7 92.0 95.3 85.9 95.0 98.0
[…]
DeiT-S (Touvron et al., 2021) De384 70.6 81.3 88.7 52.8 65.1 76.2 58.3 73.9 85.9
Hyp (Ermolov et al., 2022) De384 77.8 86.6 91.9 86.4 92.2 95.5 83.3 93.5 97.4
HIER (Kim et al., 2023) De384 78.7 86.8 92.0 88.9 93.9 96.6 83.0 93.1 97.2
LHIER (ours) De384 78.3 86.2 91.8 88.7 93.4 96.4 82.8 92.9 96.9
DINO (Caron et al., 2021) DN384 70.8 81.1 88.8 42.9 53.9 64.2 63.4 78.1 88.3
Hyp (Ermolov et al., 2022) DN384 80.9 87.6 92.4 89.2 94.1 96.7 85.1 94.4 97.8
HIER  (Kim et al., 2023) DN384 81.1 88.2 93.3 91.3 95.2 97.1 85.7 94.6 97.8
LHIER (ours) DN384 81.3 88.4 93.3 91.5 95.1 97.6 85.9 95.0 98.1
ViT-S (Dosovitskiy et al., 2021) V384 83.1 90.4 94.4 47.8 60.2 72.2 62.1 77.7 89.0
Hyp (Ermolov et al., 2022) V384 85.6 91.4 94.8 86.5 92.1 95.3 85.9 94.9 98.1
HIER  (Kim et al., 2023) V384 85.7 91.3 94.4 88.3 93.2 96.1 86.1 95.0 98.0
LHIER (ours) V384 86.2 92.1 95.2 88.6 93.6 96.3 86.7 95.3 98.3
4.2 Standard Classification Problem

### v3

@@ 3.3 Riemannian AdamW Optimizer
Algorithm 1 Curvature Learning Aware Optimization
  Given parameters $\bm{\theta}=[\bm{\theta}_{euclid},\bm{\theta}_{\mathbb{L}},\bm{\theta}_{K}]$ and gradients denoted $\mathcal{G}$
  function Optimizer Step
     for $\bm{p}$ in $\bm{\theta}_{\mathbb{L}}$ do
        Vanilla Riemannian Optimizer Step
     end for
     for $\bm{p}$ in $\bm{\theta}_{euclid}\cup\bm{\theta}_{K}$ do
        Vanilla Euclidean Optimizer Step
     end for
     MoveParameters()
  end function
  function Move Parameters
     for $\bm{p}$ in $\bm{\theta}_{\mathbb{L}}$ do
        $\bm{\mathcal{G}}_{temp}=\mathrm{PT}^{K_{t-1}}_{{\bm{p}}\rightarrow{\bm{\overline{\bm{0}}}_{t-1}}}\left({\bm{\mathcal{G}}}\right)$
        $\bm{z}=\log_{\bm{\overline{\bm{0}}}_{t-1}}^{K_{t-1}}(\bm{p})$
        $\bm{p}=\exp_{\overline{\bm{0}}_{t}}^{K_{t}}(\bm{z})$
        $\bm{\mathcal{G}}=\mathrm{PT}^{K_{t}}_{{\bm{\overline{\bm{0}}}_{t}}\rightarrow{\bm{p}}}\left({\bm{\mathcal{G}}_{temp}}\right)$
     end for
  end function
3.4 Maximum Distance Rescaling
Background
@@ 3.4 Maximum Distance Rescaling
To constrain hyperbolic vectors within a specified maximum distance, either a normalization function or a parameter clipping method is required. Parameter clipping can be challenging to train, as it may lead to information loss and introduce non-smooth gradients. On the other hand, common normalization functions like tanh and the sigmoid function tend to saturate quickly, limiting their effectiveness as seen in the sigmoid implementation by Chen et al. (2022).
Flexible Scaling Function
To address these issues, we introduce a modified scaling function, designed to provide finer control over both the maximum values and the slope of the curve. A visualization of this function is provided in Figure 2, and the formulation is presented below:
@@
where $y\in\mathbb{R}^{d}$, $m$ is our desired maximum value, and $s$ controls the slope of the curve. We now have a maximum distance value to adhere to and a flexible distance normalizing function. To apply this to the hyperbolic vectors, we suggest performing the scaling on the tangent plane of the origin. However, this is an expensive operation to perform repeatedly, as such we derive in Appendix C the equivalent factorized form for the scaling of the space values:
@@
where $D(\bm{x},\overline{\bm{0}})_{rescaled}^{k}$ is the scaled distances by plugging in $D(\bm{x},\overline{\bm{0}})$ and $D^{K}_{\overline{\bm{0}}_{max}}$ in Equation 3. The time component of the vector is recalculated based on the norm of its spatial components, providing a complete scaling operation for the Lorentz space. We apply this tanh scaling during transitions between manifolds, including from Euclidean to Lorentz spaces and between Lorentz spaces of different curvatures. It is also used after Lorentz Boosts and direct Lorentz concatenations (Qu & Zou, 2022), as well as following variance-based rescaling in the batch normalization layer, where adjustments can push points outside the valid radius, often resulting in invalid values.
4 Domain Adaption
In order to empirically prove the effectiveness of our proposed solutions, we apply them to metric learning, EEG classification, image classification, and image generation tasks. In the following, we tackle each of these tasks by offering a brief overview of the problem setting with an existing reference model, our hyperbolic adaptation for this problem, and the goal from this experimental setup.
@@ 4.1 Hierarchical Metric Learning Problem
Moving to Lorentz Space
To adapt the HIER model to the hyperboloid, we first replace the Euclidean layer norm and linear layer with a Lorentzian norm and linear layer (Bdeir et al., 2024). We then modify the HIER loss by setting the hierarchical proxies as learnable hyperbolic parameters and replace the Poincaré distance with the Lorentzain distance. This allows us to optimize the hyperbolic parameters directly on the manifold using our new optimization schema. We also apply distance rescaling before layer norm and after the linear layer. We refer to this setup as LHIER.
Experimental Goals
Kim et al. (2023) employ the Euclidean AdamW optimizer with Euclidean parameterizations of hyperbolic proxies. Their experiments use FP16 precision. As the setting is already hyperbolic, significant gains from transitioning to Lorentz space are unlikely. We use this setup to evaluate our components’ ability to learn curvature in low-precision, unstable environments. Any improvements likely stem from better curvature adaptation to the data and task.
@@
We follow the experimental setup in Kim et al. (2023) and rely on four main datasets: CUB-200-2011 (CUB)(Welinder et al., 2010), Cars-196 (Cars)(Krause et al., 2013), Stanford Online Product (SOP)(Song et al., 2016), and In-shop Clothes Retrieval (InShop)(Liu et al., 2016). Performance is measured using Recall@k, the fraction of queries with at least one relevant sample in their k-nearest neighbors. All model backbones are pre-trained on ImageNet for fair comparisons.
Results
Table 1: Performance of metric learning methods on the four datasets as provided by (Kim et al., 2023). All architecture backbones are pretrained and tested with the new LHIER loss. Superscripts denote their embedding dimensions and $\dagger$ indicates models using larger input images. As in (Kim et al., 2023), network architectures are abbreviated as, R–ResNet50 (Wang et al., 2017), B–Inception with BatchNorm (Ioffe & Szegedy, 2015), De–DeiT (Touvron et al., 2021), DN–DINO (Caron et al., 2021) and V–ViT (Dosovitskiy et al., 2020)
Methods Arch. CUB Cars SOP
R@1 R@2 R@1 R@2 R@1 R@10
CNN Backbone
NSoftmax R128 56.5 69.6 81.6 88.7 75.2 88.7
MIC R128 66.1 76.8 82.6 89.1 77.2 89.4
XBM R128 - - - - 80.6 91.6
XBM B512 65.8 75.9 82.0 88.7 79.5 90.8
HTL B512 57.1 68.8 81.4 88.0 74.8 88.3
MS B512 65.7 77.0 84.1 90.4 78.2 90.5
SoftTriple B512 65.4 76.4 84.5 90.7 78.6 86.6
PA B512 68.4 79.2 86.1 91.7 79.1 90.8
NSoftmax R512 61.3 73.9 84.2 90.4 78.2 90.6
ProxyNCA++ R512 69.0 79.8 86.5 92.5 80.7 92.0
Hyp R512 65.5 76.2 81.9 88.8 79.9 91.5
HIER R512 70.1 79.4 88.2 93.0 80.2 91.5
LHIER R512 73.4 82.4 90.0 94.0 81.9 93.1
Increase % 4.71 3.78 2.04 1.08 1.49 1.20
As shown in Table 1, LHIER learns curvature without stability issues. Our approach improves model performance, with recall@1 gains ranging from 1.49% to 4.71%. Additional ablation experiments with different optimizers and configurations are detailed in Table 7.
4.2 EEG Classification Problem
Problem Setting
Given EEG recordings labeled into distinct categories, the goal is to classify new recordings based on signal patterns. Each recording is a time series $X\in\mathbb{R}^{C\times T}$, where $C$ is the number of channels and $T$ is the time length. Each label is $y\in\{1,\dots,K\}$. Given $N$ labeled recordings $(X_{1},y_{1}),\dots,(X_{N},y_{N})$ from an unknown distribution $p$, the task is to train a model $\hat{y}$ that maps EEG signals $X$ to their correct class.
Reference Model
The reference model MAtt (Pan et al., 2022) has already taken steps towards improving data preparation. In the feature extraction phase, MAtt uses 2 convolutions along the time and sensor dimensions respectively. The embeddings are transformed into Symmetric Positive Definite (SPD) data by mapping the feature embeddings from Euclidean space onto the SPD manifold. The resulting sequence is then fed into a Manifold Attention Model and the outputs are projected back to the Euclidean space and classified using a fully connected layer.
@@ 4.2 EEG Classification Problem
To adapt Matt to a Lorentz space setting, we project the embeddings from the SPD manifold to the Lorentz manifold after the input slicing and sample covariance matrix projection. We then pass them through an adapted Lorentz Attention layer that replaces the SPD layers, distance function, and centroid with the Lorentzian counterparts. Finally, we use the Lorentz MLR (Bdeir et al., 2024) as a final classification head. We refer to this enhanced model as HyperMAtt.
Experimental Goals
While this process already represents EEG sequences in Riemannian space, we enhance it further with our Riemannian AdamW Optimizer. Hyperbolic spaces are prone to overfitting, especially in EEG classification, where models are trained per subject with limited sessions, leading to data scarcity. The proposed optimizer improves regularization, ensuring smoother and more balanced convergence. We avoid curvature learning, as it could worsen overfitting and reduce performance, but retain distance rescaling functions.
Experimental Setup
We compare our model HyperMAtt on three common EEG datasets for the domain of EEG classification. These datasets are MI (Motor Imagery) (Brunner et al., 2008), SSVEP (Steady-State Visual Evoked Potentials) (Nikolopoulos, 2021) and ERN (Error-Related Negativity) (Margaux et al., 2012) representing three different tasks, motor imagery, visual stimuli, and error recognition, respectively. Here, we follow the same preprocessing procedure and train/val/test splits as MAtt. For the datasets MI and SSVEP, we use Accuracy as an evaluation metric and AUC.
For the application domain of EEG, we evaluate our method against current state-of-the-art models for EEG Classification including Inception (Burchert et al., 2024), MAtt (Pan et al., 2022), MBEEGSE (Altuwaijri et al., 2022), FBCNet (Mane et al., 2021), TCNet-Fusion (Musallam et al., 2021), EEG-TCNet (Ingolfsson et al., 2020), SCCNet (Wei et al., 2019), EEGNet (Lawhern et al., 2018), and ShallowConvNet (Schirrmeister et al., 2017).
@@ Results
In Table 2 we show the performance of HyperMAtt compared to the current state-of-the-art baselines, where the results were aggregated from (Pan et al., 2022; Burchert et al., 2024). The performance for HyperMAtt is measured across 3 runs and we report the mean and standard deviation. Here, our new optimizer archives state-of-the-art results for SSVEP and ERN, the two comparatively smaller datasets, which are more prone to overfitting. We also show the improved performance of our RAdamW optimizer vs the existing RAdam by training the model using both and comparing the results in Table 8. We also achieve these results in significantly less time/epoch compared to MAtt. Specifically for MI 0.077s/epoch vs 3.4s/epoch, for SSVEP 0.081s/epoch vs 4s/epochs, and for ERN 0.061s/epoch vs 1.9s/epoch.
Table 2: Performance comparison for the EEG datasets MI, SSVEP, and ERN. We report the average accuracy for MI and SSVEP and the AUC for ERN. The best result is highlighted in bold.
Models MI SSVEP ERN ShallowConvNet 61.84$\pm$6.39 56.93$\pm$6.97 71.86$\pm$2.64 EEGNet 57.43$\pm$6.25 53.72$\pm$7.23 74.28$\pm$2.47 SCCNet 71.95$\pm$5.05 62.11$\pm$7.70 70.93$\pm$2.31 EEG-TCNet 67.09$\pm$4.66 55.45$\pm$7.66 77.05$\pm$2.46 TCNet-Fusion 56.52$\pm$3.07 45.00$\pm$6.45 70.46$\pm$2.94 FBCNet 71.45$\pm$4.45 53.09$\pm$5.67 60.47$\pm$3.06 MBEEGSE 64.58$\pm$6.07 56.45$\pm$7.27 75.46$\pm$2.34 Inception 62.85$\pm$3.21 62.71$\pm$2.95 73.55$\pm$5.08 MAtt 74.71$\pm$5.01 65.50$\pm$8.20 76.01$\pm$2.28 HyperMAtt (ours) 74.12$\pm$2.91 68.10$\pm$2.41 78.01$\pm$1.30 Increase in % -0.79 3.97 1.25
4.3 Standard Image Classification Problem
Problem Setting and Reference Model
In their work, Bdeir et al. (2024) proposed a fully hyperbolic 2D convolutional layer by breaking down the convolution operation into a window-unfolding step followed by a modified version of the Lorentz Linear Layer from Chen et al. (2022). This approach ensured that the convolution outputs remained on the hyperboloid. However, the manual patch creation combined with matrix multiplication made the computation extremely expensive, as it prevented the use of highly optimized CUDA implementations for convolutions. As such, the authors proposed a two versions of their hyperbolic ResNet classifier. A fully hyperbolic model with all Lorentz ResNet blocks (HCNN) and a hybrid encoder model with alternating Euclidean and hyperbolic blocks (HECNN).
Reference Optimization
To adapt our methods for both models, we introduce the distance rescaling function after every convolution in the encoder. Additionally, we attempt to improve the performance by performing a parametrization trick on existing CUDA convolutions to ensure hyperbolic outputs. This is done by parametrizing the convolution weight as a rotation operation before passing the input, and then applying a boost operation afterwards. The end result is then equivalent to the convolution operation by Bdeir et al. (2024) while mitigating computation complexity. Specific implementation details can be found in Appendix B. We denote our Hybrid and Fully hyperbolic models as HECNN+ and HCNN+ respectively.
Experimental Goals
In their paper, (Bdeir et al., 2024) notice lower performance in the fully hyperbolic models compared to the hybrid models and attribute this to instability in training. Additionally, they cite crashes and model divergence when learning the curvature instead of setting it to constant. This would then be an ideal setting to test the optimization schema and rescaling function and their impact on performance and stability. We do not use the RAdamW optimizer in this setting as RSGD is already used and employs proper L2-regularization.